OpenStudy (anonymous):

Which expressions are equivalent to -40+24i? Choose exactly two answers that are correct. A. 4i(6-10i) B. 4i(6+10i) C. 8i(3-5i) D. 8i(3+5i)

Nnesha (nnesha):

apply distributive property distribute parentheses by outside number/variable \[\huge\rm \color{reD}{a}(b+c) =\color{red}{a} ·b +\color{reD}{a}·c =\color{ReD}{a}b +\color{ReD}{a}c\]

OpenStudy (anonymous):

Uh, can you help me through it? Sorry.

Nnesha (nnesha):

alright let's do a example \[\huge\rm 3i(2+5i)\] multiply both terms in the parentheses by outside number/variable which is 3i so multiply 3i times 2 + 3i times 5i\[\huge\rm 3i·2+3i·5i\] \[\large\rm (3·2)i+(3·5)i·i\]when we multiply same varaibles we should add their exponents so i^1 · i^1 =i^{1+1} =i^2 \[\large\rm (3·2)i+(3·5)i^2\] and i ^2 =-1 so you can replace i^2 with -1\[\rm 6i+15(-1) =6i-15\]

Nnesha (nnesha):

so A: 4i(6-10i) distribute: multiply both terms in the parentheses by 4i